OpenStudy (anonymous):

the point (11,12) is on the terminal side of an angle in standard position. Determine the exact value of sinx. Please show how you got the answer!!

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sin(\theta)=\frac{12}{\sqrt{11^2+12^2}}\] as the hypotenuse of that right triangle is \[\sqrt{11^2+12^2}\] by pythagoras

OpenStudy (anonymous):

how do you know where the numbers go on the triangle?